# Discrete Neural Flow Samplers with Locally Equivariant Transformer

### v1

@@ 4 Applications and Experiments
To support our theoretical discussion, we first evaluate the proposed methods by sampling from predefined unnormalised distributions. We then demonstrate two important applications to DNFS: i) training discrete energy-based models and ii) solving combinatorial optimisation problems. Detailed experimental settings and additional results are provided in Appendix E.
@@ 4.1 Sampling from Unnormalised Distributions
The results, shown in Figure 4, demonstrate that the proposed method, DNFS, produces samples that closely resemble those from the oracle Gibbs sampler. In contrast, GFlowNet occasionally suffers from mode collapse, particularly on structured datasets such as the checkerboard pattern. Although LEAPS with leTF achieves performance comparable to DNFS, it sometimes produces inaccurate samples that fall in smoother regions of the energy landscape, potentially due to imprecise estimation of $\partial_{t}\log Z_{t}$. Furthermore, we observe that LEAPS with leConv performs poorly in this setting (see Figure 12), reinforcing the limited expressiveness of locally equivariant convolutional networks when applied to non-grid data structures. For a more comprehensive evaluation, additional visualisations on other datasets are provided in Figure 11, further illustrating the effectiveness of our method.
@@
where $J=\sigma A_{D}$ with $\sigma\in\mathbb{R}$ and $A_{D}$ being the adjacency matrix of a $D\times D$ grid. [^2] The adjacency matrix is constructed using A_D = igraph.Graph.Lattice(dim=[D, D], circular=True). In Figure 5, we evaluate DNFS on a $D=10\times 10$ lattice grid with $\sigma=0.1$, comparing it to baselines methods in terms of effective sample size (ESS) (see Section D.1 for details) and the energy histogram of $5,000$ samples. The oracle energy distribution is approximated using long-run Gibbs sampling. The results show that DNFS performs competitively with LEAPS and significantly outperforms GFlowNet, which fails to capture the correct mode of the energy distribution. Although LEAPS with leConv achieves a comparable effective sample size, it yields a less accurate approximation of the energy distribution compared to DNFS. Furthermore, DNFS attains a lower loss value, as shown in Figure 15.
@@
These findings underscore the effectiveness of the proposed neural sampler learning algorithm and highlight the strong generalisation capability of leTF across both grid-structured and non-grid data. For a more comprehensive evaluation, we also compare our method with MCMC-based approaches in Figure 13, further demonstrating the efficacy of our approach.
@@ 4.2 Training Discrete Energy-based Models
To approximate this intractable gradient, we train a rate matrix $R_{t}^{\theta}$ to sample from the target $p_{\phi}$. This enables using importance sampling to estimate the expectation (see Section D.1 for details).
@@
Figure 6 illustrates the estimated energy landscape alongside samples generated using the trained DNFS sampler. The results demonstrate that the learned EBM effectively captures the multi-modal structure of the underlying distribution, accurately modelling the energy across the data support. The sampler produces samples that closely resemble the training data, highlighting the effectiveness of DNFS in training discrete EBMs. Additional qualitative results are presented in Figure 16. In Table 2, we provide a quantitative comparison of our method with several baselines, focusing in particular on two contrastive divergence (CD)-based approaches: PCD (Tieleman, 2008) with MCMC and ED-GFN (Zhang et al., 2022a) with GFlowNet. Our method, built upon the proposed DNFS, consistently outperforms PCD in most settings, underscoring the effectiveness of DNFS in training energy-based models. While ED-GFN achieves better performance than DNFS, it benefits from incorporating a Metropolis-Hastings (MH) (Hastings, 1970) correction to sample from the model distribution $p_{\phi}$ in Equation 12, which may offer an advantage over the importance sampling strategy used in Equation 13. We leave the integration of MH into DNFS as a direction for future work.
@@
Training Ising Models. We further assess DNFS for training the lattice model defined in Equation 11. Following Grathwohl et al. (2021), we generate training data using Gibbs sampling and use these samples to learn a symmetric matrix $J_{\phi}$ to estimate the true matrix in the Ising model. Importantly, the training algorithms do not have access to the true data-generating matrix $J$, but only to the synthesised samples.
In Figure 7, we consider a $D=10\times 10$ grid with $\sigma=0.1$ and visualise the learned matrix $J_{\phi}$ using a heatmap. The results show that the proposed method successfully captures the underlying pattern of the ground truth, demonstrating the effectiveness of DNFS. Further quantitative analysis across various configurations of $D$ and $\sigma$ is presented in Table 4.
@@ 4.3 Solving Combinatorial Optimisation Problems
Therefore, we can train DNFS to sample from $p(x)$, which will produce high-quality solutions to the MIS problem. To enable generalisation across different graphs $G$, we condition the locally equivariant transformer on the graph structure. Specifically, we incorporate the Graphformer architecture (Ying et al., 2021), which adjusts attention weights based on the input graph. This allows the model to adapt to varying graph topologies. We refer to this architecture as the locally equivariant Graphformer (leGF), with implementation details provided in Section D.4.
Experimental Settings. In this experiment, we apply our method to solve the Maximum Independent Set (MIS) problem, with other settings deferred to Section E.3. Specifically, we benchmark MIS on Erdős–Rényi (ER) random graphs (Erdos, 1961), comprising 1,000 training and 100 testing instances, each with 16 to 75 vertices. Evaluation on the test set includes both performance and inference time. We report the average solution size and the approximation ratio with respect to the best-performing mixed-integer programming solver (Gurobi) (Gurobi, 2023), which serves as the oracle.
Results & analysis. We compare our method against two baselines: an annealed MCMC sampler (Sun et al., 2023b) using DMALA (Zhang et al., 2022b), and a neural sampler based on GFlowNet (Zhang et al., 2023a). Additionally, we include results from a randomly initialised version of DNFS without training, which serves as an estimate of the task’s intrinsic difficulty. As shown in Table 1, DNFS after training substantially outperforms its untrained counterpart, highlighting the effectiveness of our approach. While the MCMC-based method achieves the strongest overall performance, it requires longer inference time. Compared to GFlowNet, another neural sampler, DNFS performs slightly worse. This may be attributed to the fact that GFlowNet restricts sampling to feasible solutions only along the trajectory, effectively reducing the exploration space and making the learning problem easier. Incorporating this inductive bias into DNFS is a promising direction for future work. Nevertheless, a key advantage of our method is that the unnormalised marginal distribution $p_{t}$ is known, allowing us to integrate additional MCMC steps to refine the sampling trajectory. As shown in the last row of Table 1, this enhancement leads to a substantial performance gain. Further analysis of this approach is provided in Table 6 in the appendix.
@@ 5 Related Work
CTMCs and Discrete Diffusion. Our work builds on the framework of continuous-time Markov chains (CTMCs), which were first introduced in generative modelling by Austin et al. (2021); Sun et al. (2023c); Campbell et al. (2022) under the context of continuous-time discrete diffusion models, where the rate matrix is learned from training data. This approach was later simplified and generalised to discrete-time masked diffusion (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2024), demonstrating strong performance across a wide range of applications, including language modelling (Lou et al., 2024), molecular simulation (Campbell et al., 2024), and code generation (Gat et al., 2024). However, these methods require training data and are inapplicable when only an unnormalised target is given.
@@
MCMC and Neural Samplers. Markov chain Monte Carlo (MCMC) (Metropolis et al., 1953) is the de facto approach to sampling from a target distribution. In discrete spaces, Gibbs sampling (Casella & George, 1992) is a widely adopted method. Building on this foundation, Zanella (2020) improve the standard Gibbs method by incorporating locally informed proposals to improve sampling efficiency. This method was extended to include gradient information to drastically reduce the computational complexity of flipping bits in several places. This idea was further extended by leveraging gradient information (Grathwohl et al., 2021; Sun et al., 2022a), significantly reducing the computational cost. Inspired by these developments, discrete analogues of Langevin dynamics have also been introduced to enable more effective sampling in high-dimensional discrete spaces (Zhang et al., 2022b; Sun et al., 2023a). Despite their theoretical appeal, MCMC methods often suffer from slow mixing and poor convergence in practice. To address these limitations, recent work has proposed neural samplers, including diffusion-based (Vargas et al., 2024; Chen et al., 2024; Richter & Berner, 2024) and flow-based (Máté & Fleuret, 2023; Tian et al., 2024; Chen et al., 2025) approaches. However, the majority of these methods are designed for continuous spaces, and there remains a notable gap in the literature when it comes to sampling methods for discrete distributions. A few exceptions include Sanokowski et al. (2024, 2025), which are inspired by discrete diffusion models and primarily target combinatorial optimisation problems. LEAPS (Holderrieth et al., 2025) and our method DNFS are more closely related to discrete flow models (Campbell et al., 2024; Gat et al., 2024), as both can be view as learning a CTMC to satisfy the Kolmogorov forward equation. While LEAPS parametrise $\partial_{t}\log Z_{t}$ using a neural network, DNFS estimates it via coordinate descent.
@@
Discrete EBMs and Neural Combinatorial Optimisation. Contrastive divergence is the de facto approach to train energy-based models, but it relies on sufficiently fast mixing of Markov chains, which typically cannot be achieved (Nijkamp et al., 2020). To address this, several sampling-free alternatives have been proposed, including energy discrepancy (Schröder et al., 2024), ratio matching (Lyu, 2012), and variational approaches (Lázaro-Gredilla et al., 2021). More recently, Zhang et al. (2022a) replace MCMC with GFlowNet, a neural sampler that arguably offers improvement by reducing the risk of getting trapped in local modes. Our work follows this line of research by using DNFS as a neural alternative to MCMC for training energy-based models. Sampling methods are also widely used to solve combinatorial optimisation problems (COPs). Early work (Sun et al., 2022b) demonstrated the effectiveness of MCMC techniques for this purpose. More recent approaches (Zhang et al., 2023a; Sanokowski et al., 2024, 2025) leverage neural samplers to learn amortised solvers for COPs. In this paper, we further show that DNFS is well-suited for combinatorial optimisation tasks, demonstrating its flexibility and broad applicability.
@@ 6 Conclusion and Limitation
We proposed discrete neural flow samplers (DNFS), a discrete sampler that learns a continuous-time Markov chain to satisfy the Kolmogorov forward equation. While our empirical studies demonstrate the effectiveness of DNFS across various applications, it also presents several limitations. A natural direction for future work is to extend DNFS beyond binary settings. However, this poses significant challenges due to the high computational cost of evaluating the ratios in Equation 7. As demonstrated in Section C.2, a naive Taylor approximation introduces bias into the objective, resulting in suboptimal solutions. Overcoming this limitation will require more advanced and principled approximation techniques. Additionally, we find that the current framework struggles to scale to very high-dimensional distributions. This difficulty arises mainly from the summation over the ratios in Equation 7, which can lead to exploding loss values. Designing methods to stabilise this computation represents a promising avenue for future research. Finally, extending DNFS to the masked diffusion setting offers another compelling direction, with the potential to support more flexible and efficient sampling over structured discrete spaces.
@@ Appendix A Variance Reduction and Control Variates
In this section, we analyse the variance reduction phenomenon shown in Figure 1 through the lens of control variates. We begin with the proof of Equation 8 and the discussion of control variates, then establish its connection to LEAPS (Holderrieth et al., 2025).
A.1 Proof of Equation 8
Recall from Equation 8 that the following equation holds:
@@ A.1 Proof of Equation 8
We are now ready to prove Equation 8. Specifically:
@@
where the second and third equations follow Lemmas 2 and 1 respectively.
@@ A.2 Discrete Stein Control Variates
To better understand the role of control variates (Geffner & Domke, 2018) in variance reduction as shown in Figure 1, let us consider a standard Monte Carlo estimation problem. Suppose our goal is to estimate the expectation $\mu={\mathbb{E}}_{\pi}[f(x)]$, where $f(x)$ is a function of interest. A basic estimator for $\mu$ is the the Monte Carlo average $\hat{\mu}=\frac{1}{K}\sum_{k=1}^{K}f(x^{(k)}),x^{(k)}\sim\pi$. Now, suppose we have access to another function $g(x)$, known as a control variate, which has a known expected value $\gamma={\mathbb{E}}_{\pi}[g(x)]$. We can the use $g(x)$ to construct a new estimator: $\check{\mu}=\frac{1}{K}\sum_{k=1}^{K}(f(x^{(k)})-\beta g(x^{(k)}))+\beta\gamma$. This new estimator $\check{\mu}$ is unbiased for any choice of $\beta$, since ${\mathbb{E}}[\check{\mu}]={\mathbb{E}}[f(x)]-\beta{\mathbb{E}}[g(x)]+\beta\gamma=\mu$. The benefit of this construction lies not in bias correction but in variance reduction. To see this, we can compute the variance of $\check{\mu}$:
@@
Fortunately, Lemma 1 provides a principled way to construct a control variate tailored to our setting ${\mathbb{E}}_{p_{t}}[f(x)]\triangleq{\mathbb{E}}_{p_{t}}[\partial_{t}\log\tilde{p}_{t}(x)]\approx\frac{1}{K}\sum_{k=1}^{K}\partial_{t}\log\tilde{p}_{t}(x^{(k)})$, where $x^{(k)}\sim p_{t}$. To reduce the variance of this estimator, we seek a control variate $g(x)$ whose expectation under $p_{t}$ is known. Inspired by the discrete Stein identity, we define $g(x)=\sum_{y}R_{t}(x,y)\frac{p_{t}(y)}{p_{t}(x)}$, which satisfies ${\mathbb{E}}_{p_{t}}[g(x)]=0$ by Lemma 1. This makes $g$ a valid control variate with known mean. Using this construction, we can define a variance-reduced estimator as:
@@
This estimator remains unbiased for any $\beta$, but with the optimal choice $\beta^{*}$, it can substantially reduce variance. Moreover, in the special case where the parameter $\theta$ is optimal (in the sense that the objective in Equation 5 equals zero), an even stronger result emerges: the control variate becomes perfectly (negatively) correlated with the target function. That is, $g(x)=-f(x)+c$, where $c$ is a constant independent of the sample $x$, leading to $\mathrm{Corr}(f,g)=-1$. In this idealised case, $\check{\mu}$ becomes a zero-variance estimator, a rare but highly desirable scenario. For a more comprehensive discussion of discrete Stein-based control variates and their applications in variance reduction, we refer the reader to Shi et al. (2022).

### v2

@@ 4 Applications and Experiments
To support our theoretical discussion, we first evaluate the proposed methods by sampling from predefined unnormalised distributions. We then demonstrate two important applications to DNFS: i) training discrete energy-based models and ii) solving combinatorial optimisation problems. Detailed experimental settings and additional results are provided in Appendix˜E.
@@ 4.1 Sampling from Unnormalised Distributions
The results, shown in Figure˜4, demonstrate that the proposed method, DNFS, produces samples that closely resemble those from the oracle Gibbs sampler. In contrast, GFlowNet occasionally suffers from mode collapse, particularly on structured datasets such as the checkerboard pattern. Although LEAPS with leTF achieves performance comparable to DNFS, it sometimes produces inaccurate samples that fall in smoother regions of the energy landscape, potentially due to imprecise estimation of $\partial_{t}\log Z_{t}$. Furthermore, we observe that LEAPS with leConv performs poorly in this setting (see Figure˜12), reinforcing the limited expressiveness of locally equivariant convolutional networks when applied to non-grid data structures. For a more comprehensive evaluation, additional visualisations on other datasets are provided in Figure˜11, further illustrating the effectiveness of our method.
@@
where $J=\sigma A_{D}$ with $\sigma\in\mathbb{R}$ and $A_{D}$ being the adjacency matrix of a $D\times D$ grid. [^2] The adjacency matrix is constructed using A_D = igraph.Graph.Lattice(dim=[D, D], circular=True). In Figure˜5, we evaluate DNFS on a $D=10\times 10$ lattice grid with $\sigma=0.1$, comparing it to baselines methods in terms of effective sample size (ESS) (see Section˜D.1 for details) and the energy histogram of $5,000$ samples. The oracle energy distribution is approximated using long-run Gibbs sampling. The results show that DNFS performs competitively with LEAPS and significantly outperforms GFlowNet, which fails to capture the correct mode of the energy distribution. Although LEAPS with leConv achieves a comparable effective sample size, it yields a less accurate approximation of the energy distribution compared to DNFS. Furthermore, DNFS attains a lower loss value, as shown in Figure˜15.
@@
These findings underscore the effectiveness of the proposed neural sampler learning algorithm and highlight the strong generalisation capability of leTF across both grid-structured and non-grid data. For a more comprehensive evaluation, we also compare our method with MCMC-based approaches in Figure˜13, further demonstrating the efficacy of our approach.
@@ 4.2 Training Discrete Energy-based Models
To approximate this intractable gradient, we train a rate matrix $R_{t}^{\theta}$ to sample from the target $p_{\phi}$. This enables using importance sampling to estimate the expectation (see Section˜D.1 for details).
@@
Figure˜6 illustrates the estimated energy landscape alongside samples generated using the trained DNFS sampler. The results demonstrate that the learned EBM effectively captures the multi-modal structure of the underlying distribution, accurately modelling the energy across the data support. The sampler produces samples that closely resemble the training data, highlighting the effectiveness of DNFS in training discrete EBMs. Additional qualitative results are presented in Figure˜16. In Table˜3, we provide a quantitative comparison of our method with several baselines, focusing in particular on two contrastive divergence (CD)-based approaches: PCD (Tieleman, 2008) with MCMC and ED-GFN (Zhang et al., 2022a) with GFlowNet. Our method, built upon the proposed DNFS, consistently outperforms PCD in most settings, underscoring the effectiveness of DNFS in training energy-based models. While ED-GFN achieves better performance than DNFS, it benefits from incorporating a Metropolis-Hastings (MH) (Hastings, 1970) correction to sample from the model distribution $p_{\phi}$ in Equation˜12, which may offer an advantage over the importance sampling strategy used in Equation˜13. We leave the integration of MH into DNFS as a direction for future work.
@@
Training Ising Models. We further assess DNFS for training the lattice model defined in Equation˜11. Following Grathwohl et al. (2021), we generate training data using Gibbs sampling and use these samples to learn a symmetric matrix $J_{\phi}$ to estimate the true matrix in the Ising model. Importantly, the training algorithms do not have access to the true data-generating matrix $J$, but only to the synthesised samples.
In Figure˜7, we consider a $D=10\times 10$ grid with $\sigma=0.1$ and visualise the learned matrix $J_{\phi}$ using a heatmap. The results show that the proposed method successfully captures the underlying pattern of the ground truth, demonstrating the effectiveness of DNFS. Further quantitative analysis across various configurations of $D$ and $\sigma$ is presented in Table˜5.
@@ 4.3 Solving Combinatorial Optimisation Problems
Therefore, we can train DNFS to sample from $p(x)$, which will produce high-quality solutions to the MIS problem. To enable generalisation across different graphs $G$, we condition the locally equivariant transformer on the graph structure. Specifically, we incorporate the Graphformer architecture (Ying et al., 2021), which adjusts attention weights based on the input graph. This allows the model to adapt to varying graph topologies. We refer to this architecture as the locally equivariant Graphformer (leGF), with implementation details provided in Section˜D.4.
Experimental Settings. In this experiment, we apply our method to solve the Maximum Independent Set (MIS) problem, with other settings deferred to Section˜E.3. Specifically, we benchmark MIS on Erdős–Rényi (ER) random graphs (Erdos, 1961), comprising 1,000 training and 100 testing instances, each with 16 to 75 vertices. Evaluation on the test set includes both performance and inference time. We report the average solution size and the approximation ratio with respect to the best-performing mixed-integer programming solver (Gurobi) (Gurobi, 2023), which serves as the oracle.
Results & analysis. We compare our method against two baselines: an annealed MCMC sampler (Sun et al., 2023b) using DMALA (Zhang et al., 2022b), and a neural sampler based on GFlowNet (Zhang et al., 2023a). Additionally, we include results from a randomly initialised version of DNFS without training, which serves as an estimate of the task’s intrinsic difficulty. As shown in Table˜1, DNFS after training substantially outperforms its untrained counterpart, highlighting the effectiveness of our approach. While the MCMC-based method achieves the strongest overall performance, it requires longer inference time. Compared to GFlowNet, another neural sampler, DNFS performs slightly worse. This may be attributed to the fact that GFlowNet restricts sampling to feasible solutions only along the trajectory, effectively reducing the exploration space and making the learning problem easier. Incorporating this inductive bias into DNFS is a promising direction for future work. Nevertheless, a key advantage of our method is that the unnormalised marginal distribution $p_{t}$ is known, allowing us to integrate additional MCMC steps to refine the sampling trajectory. As shown in the last row of Table˜1, this enhancement leads to a substantial performance gain. Further analysis of this approach is provided in Table˜7 in the appendix.
@@ 5 Related Work
CTMCs and Discrete Diffusion. Our work builds on the framework of continuous-time Markov chains (CTMCs), which were first introduced in generative modelling by Austin et al. (2021); Sun et al. (2023c); Campbell et al. (2022) under the context of continuous-time discrete diffusion models, where the rate matrix is learned from training data. This approach was later simplified and generalised to discrete-time masked diffusion (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2024), demonstrating strong performance across a wide range of applications, including language modelling (Lou et al., 2024; Zhang et al., 2025), molecular simulation (Campbell et al., 2024), and code generation (Gat et al., 2024; Gong et al., 2025). However, these methods require training data and are inapplicable when only an unnormalised target is given.
@@
MCMC and Neural Samplers. Markov chain Monte Carlo (MCMC) (Metropolis et al., 1953) is the de facto approach to sampling from a target distribution. In discrete spaces, Gibbs sampling (Casella & George, 1992) is a widely adopted method. Building on this foundation, Zanella (2020) improve the standard Gibbs method by incorporating locally informed proposals to improve sampling efficiency. This method was extended to include gradient information to drastically reduce the computational complexity of flipping bits in several places. This idea was further extended by leveraging gradient information (Grathwohl et al., 2021; Sun et al., 2022a), significantly reducing the computational cost. Inspired by these developments, discrete analogues of Langevin dynamics have also been introduced to enable more effective sampling in high-dimensional discrete spaces (Zhang et al., 2022b; Sun et al., 2023a). Despite their theoretical appeal, MCMC methods often suffer from slow mixing and poor convergence in practice. To address these limitations, recent work has proposed neural samplers, including diffusion-based (Vargas et al., 2024; Chen et al., 2024; Richter & Berner, 2024) and flow-based (Máté & Fleuret, 2023; Tian et al., 2024; Chen et al., 2025) approaches. However, the majority of these methods are designed for continuous spaces, and there remains a notable gap in the literature when it comes to sampling methods for discrete distributions. A few exceptions include Sanokowski et al. (2024, 2025), which are inspired by discrete diffusion models and primarily target combinatorial optimisation problems. A concurrent work, MDNS (Zhu et al., 2025), introduces a masked diffusion neural sampler grounded in stochastic optimal control theory (Berner et al., 2022). LEAPS (Holderrieth et al., 2025) and our method DNFS are more closely related to discrete flow models (Campbell et al., 2024; Gat et al., 2024), as both can be view as learning a CTMC to satisfy the Kolmogorov forward equation. While LEAPS parametrise $\partial_{t}\log Z_{t}$ using a neural network, DNFS estimates it via coordinate descent.
@@
Discrete EBMs and Neural Combinatorial Optimisation. Contrastive divergence is the de facto approach to train energy-based models, but it relies on sufficiently fast mixing of Markov chains, which typically cannot be achieved (Nijkamp et al., 2020). To address this, several sampling-free alternatives have been proposed, including energy discrepancy (Schröder et al., 2023, 2024), ratio matching (Lyu, 2012), and variational approaches (Lázaro-Gredilla et al., 2021). More recently, Zhang et al. (2022a) replace MCMC with GFlowNet, a neural sampler that arguably offers improvement by reducing the risk of getting trapped in local modes. Our work follows this line of research by using DNFS as a neural alternative to MCMC for training energy-based models. Sampling methods are also widely used to solve combinatorial optimisation problems (COPs). Early work (Sun et al., 2022b) demonstrated the effectiveness of MCMC techniques for this purpose. More recent approaches (Zhang et al., 2023a; Sanokowski et al., 2024, 2025) leverage neural samplers to learn amortised solvers for COPs. In this paper, we further show that DNFS is well-suited for combinatorial optimisation tasks, demonstrating its flexibility and broad applicability.
@@ 6 Conclusion and Limitation
We proposed discrete neural flow samplers (DNFS), a discrete sampler that learns a continuous-time Markov chain to satisfy the Kolmogorov forward equation. While our empirical studies demonstrate the effectiveness of DNFS across various applications, it also presents several limitations. A natural direction for future work is to extend DNFS beyond binary settings. However, this poses significant challenges due to the high computational cost of evaluating the ratios in Equation˜7. As demonstrated in Section˜C.2, a naive Taylor approximation introduces bias into the objective, resulting in suboptimal solutions. Overcoming this limitation will require more advanced and principled approximation techniques. Additionally, we find that the current framework struggles to scale to very high-dimensional distributions. This difficulty arises mainly from the summation over the ratios in Equation˜7, which can lead to exploding loss values. Designing methods to stabilise this computation represents a promising avenue for future research. Finally, extending DNFS to the masked diffusion setting offers another compelling direction, with the potential to support more flexible and efficient sampling over structured discrete spaces.
Broader impact. This paper aims to advance machine learning research. While there may be potential societal impacts, none require specific mention at this time.
Acknowledgements
ZO is supported by the Lee Family Scholarship. We would like to thank Tobias Schröder and Sangwoong Yoon for their valuable discussions on the early draft of this work. ZO also thanks Peter Holderrieth for generously sharing his insights on LEAPS during their discussion at ICLR 2025.
@@ Appendix A Variance Reduction and Control Variates
In this section, we analyse the variance reduction phenomenon shown in Figure˜1 through the lens of control variates. We begin with the proof of Equation˜8 and the discussion of control variates, then establish its connection to LEAPS (Holderrieth et al., 2025).
A.1 Proof of Equation˜8
Recall from Equation˜8 that the following equation holds:
@@ A.1 Proof of Equation˜8
We are now ready to prove Equation˜8. Specifically:
@@
where the second and third equations follow Lemmas˜2 and 1 respectively.
@@ A.2 Discrete Stein Control Variates
To better understand the role of control variates (Geffner & Domke, 2018) in variance reduction as shown in Figure˜1, let us consider a standard Monte Carlo estimation problem. Suppose our goal is to estimate the expectation $\mu={\mathbb{E}}_{\pi}[f(x)]$, where $f(x)$ is a function of interest. A basic estimator for $\mu$ is the the Monte Carlo average $\hat{\mu}=\frac{1}{K}\sum_{k=1}^{K}f(x^{(k)}),x^{(k)}\sim\pi$. Now, suppose we have access to another function $g(x)$, known as a control variate, which has a known expected value $\gamma={\mathbb{E}}_{\pi}[g(x)]$. We can the use $g(x)$ to construct a new estimator: $\check{\mu}=\frac{1}{K}\sum_{k=1}^{K}(f(x^{(k)})-\beta g(x^{(k)}))+\beta\gamma$. This new estimator $\check{\mu}$ is unbiased for any choice of $\beta$, since ${\mathbb{E}}[\check{\mu}]={\mathbb{E}}[f(x)]-\beta{\mathbb{E}}[g(x)]+\beta\gamma=\mu$. The benefit of this construction lies not in bias correction but in variance reduction. To see this, we can compute the variance of $\check{\mu}$:
@@
Fortunately, Lemma˜1 provides a principled way to construct a control variate tailored to our setting ${\mathbb{E}}_{p_{t}}[f(x)]\triangleq{\mathbb{E}}_{p_{t}}[\partial_{t}\log\tilde{p}_{t}(x)]\approx\frac{1}{K}\sum_{k=1}^{K}\partial_{t}\log\tilde{p}_{t}(x^{(k)})$, where $x^{(k)}\sim p_{t}$. To reduce the variance of this estimator, we seek a control variate $g(x)$ whose expectation under $p_{t}$ is known. Inspired by the discrete Stein identity, we define $g(x)=\sum_{y}R_{t}(x,y)\frac{p_{t}(y)}{p_{t}(x)}$, which satisfies ${\mathbb{E}}_{p_{t}}[g(x)]=0$ by Lemma˜1. This makes $g$ a valid control variate with known mean. Using this construction, we can define a variance-reduced estimator as:
@@
This estimator remains unbiased for any $\beta$, but with the optimal choice $\beta^{*}$, it can substantially reduce variance. Moreover, in the special case where the parameter $\theta$ is optimal (in the sense that the objective in Equation˜5 equals zero), an even stronger result emerges: the control variate becomes perfectly (negatively) correlated with the target function. That is, g(x)=f(x)+cg(x)=-f(x)+c, where $c$ is a constant independent of the sample $x$, leading to $\mathrm{Corr}(f,g)=-1$. In this idealised case, $\check{\mu}$ becomes a zero-variance estimator, a rare but highly desirable scenario. For a more comprehensive discussion of discrete Stein-based control variates and their applications in variance reduction, we refer the reader to Shi et al. (2022).